B is between A and C.
AB=2x-4
BC=7x+5
AC=82
What is the value of x? Show your work for credit

Answer by checkley77(12569)

AB=2x-4
BC=7x+5
AC=82
******************
2X-4+7X+5=82
9X=82+4-5
9X=81
X=81/9
X=9 ANSWER.
PROOF:
2*9-4+7*9+5=82
18-4+63+5=82
86-4=82
82=82